OpenStudy (awy):

Given the following point, (0,-18) , in rectangular coordinates. Find another point in polar coordinates with the following conditions Conditions: 1. r > 0 0<= theta < 2pi 2. r<0 0<= theta < 2pi 3. r>0 -2pi <= theta <0 4. r<0 -2pi <= theta <0

OpenStudy (accessdenied):

Did you have any ideas on where to start with this problem? I assume the problem wishes for us to convert (0, -18) into polar coordinates in these four ways.

OpenStudy (awy):

someone is here to help, amazing and yea i think i almost have this figured out

OpenStudy (awy):

from cart coords to polar coords i used r^2=x^2+y^2 and tantheta=y/x, xcannot = 0

OpenStudy (awy):

so i know that r=18 and theta = DNE and the first condition is that r > 0 0<= theta < 2pi the polar coord should be (18,3pi/2)

OpenStudy (accessdenied):

That looks good to me.

OpenStudy (awy):

so, up till the 4th condition i'm having a slight trouble.

OpenStudy (awy):

i know the polar coord is double negative and r =-18

OpenStudy (awy):

is it -pi-(3pi/2)=-pi/2 ?

OpenStudy (awy):

oh, wait i think i figured it out, its -3pi/2

OpenStudy (accessdenied):

I'm thinking if you plotted (18, -pi/2) you'd have the right point but then switching to -18 would be opposite. Yeah.

OpenStudy (awy):

the 4th condition is basically a reflection of the 3rd condition

OpenStudy (awy):

alright ty for the help.

OpenStudy (accessdenied):

No problem! :p